OpenStudy (hba):

If x and y are distinct prime numbers less than 10,which of the following cannot be the sum of x and y ?

OpenStudy (hba):

So, \[x<10 / and/y<10\]

OpenStudy (hba):

The options are 5,6,7,8

OpenStudy (hba):

So we have to see which one of them cannot be equal to x+y

OpenStudy (hba):

@waterineyes

OpenStudy (hba):

@thinker1 ?

OpenStudy (hba):

Oh so they are prime number =2,3,5,7,9

OpenStudy (hba):

I think the answer is 8

OpenStudy (anonymous):

9 is not a prime number..

OpenStudy (hba):

sorry.

OpenStudy (anonymous):

Add any of the two prime numbers..

OpenStudy (hba):

ok i have tried adding them but none of them sums up 8

OpenStudy (hba):

No my bad

OpenStudy (anonymous):

2 + 3 = 3 + 5 = 2 + 5 =

OpenStudy (hba):

distinct means diff right ?

OpenStudy (hba):

SO 6 IS MY FINAL ANSWER ^

OpenStudy (anonymous):

Here, you are getting 5, 8 and 7.. Are you getting 6 by adding any two distinct prime numbers?? Yes. distinct means Different ( Not same).. In Hindi, "Alag"..

OpenStudy (anonymous):

Are you sure it is 6?? Be careful..

OpenStudy (hba):

I don't speak hindi lol.

OpenStudy (anonymous):

Ha ha ha ha... I was just in a mood of kidding..

OpenStudy (hba):

@waterineyes yes,pretty sure.

OpenStudy (hba):

@waterineyes k.

OpenStudy (anonymous):

If you are pretty sure, then you are pretty right too...

OpenStudy (hba):

Thanks :D

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (hba):

Sorry my internet connection was not working :(

OpenStudy (anonymous):

It is okay, You need not be sorry..

OpenStudy (hba):

k,thanks.

OpenStudy (anonymous):

Great..